Fermi National Laboratory


Accelerator Update - May 21-May 23

Last week the Fermilab accelerators broke several records. Fermilab set a new integrated luminosity record with a level of 9159 nb-1. The Main Injector (MI) broke a record that wasn't reported, that of a high intensity of protons, 5.45E12, being sent to the Antiproton Source (Pbar) target for a single pulse.

Wednesday May 21
The day shift began with Tevatron (TeV) experts conducting studies, with the Antiproton Source (Pbar) stacking, and with MiniBooNe taking beam.

At 11:08 AM, the TeV suffered a sector F3 quench during studies.

The Pbar Lithium Lens tripped off at 12:51 PM on a ground fault indication. Operators contacted experts. This halted stacking. Operations rested an aging Linac RF station (LRF5), since it wasn't needed.

Operations made the Collider Detector Facility (CDF) and D-zero (D0) collision halls ready for access. The access began at 1:33 PM. Power supply technicians began work on a TeV device (T:HE17).

Switchyard and Main Injector (MI) experts began beam studies at 2:40 PM.

At 2:58 PM, Operations established beam to MiniBooNe.

The power supply techs found the reason for our many HE17 quench trips. It was due to a barrel spice failure.

Both CDF and D0 were out of access by 4 PM.

Both Switchyard and MI experts completed their studies at 5 PM.

TeV experts resumed their studies at 6 PM.

A tech completed work on a sector B1 air compressor at 10 PM.

Thursday May 22
The midnight shift began with Pbar experts continuing with their Lithium Lens replacement, with MI experts conducting A1 transfer line emittance growth studies, and with MiniBooNe taking beam.

At 3:53 AM, Pbar experts successfully tested their newly installed Lithium Lens. The following picture shows the Lens through a lead window.

Operations put the TeV through a dry squeeze at 4 AM.

At 4:32 AM, a Booster power supply (EAPS) began causing problems. Then, at 4:55 AM, a Booster RF station (BRF5) started sparking, which was why EAPS was down. Operators called in an expert who bypassed BRF5. He had EAPS operating by 5:45 AM.

Operations put the TeV into shot setup at 6:10 AM.

At 8:40 AM, Operations established store 2581 with an initial luminosity of 6.25E30. As you can see there were lots of problems with this store. Experts began their investigations. Pbar resumed stacking.

Vacuum tech began working on Booster RF vacuum at 11:20 AM. At 11:31 AM, they completed their work and Pbar resumed stacking and MiniBooNe resumed taking beam.

At 11:34, Pbar experts completed their final adjustments on the Lithium Lens.

A Pbar expert conducted 30 minutes of stacking studies at 1 PM.

A Linac expert made adjustments on LRF5 at 1:38 PM to help extend its power amplifier's life.

At 2 PM, Operations halted Pbar stacking while a Booster tech brought BRF4 back on line. During this time a Linac expert switched the Preaccelerator to operate from the I- Source to the H- Source. Stacking resumed at 2:20 PM.

Operations terminated store 2581 at 3:16 PM.

TeV expert began another round of studies at 3:53 PM.

At 7:35, a Booster power supply (B:V803) tripped off due to a high transformer temperature. Operators called in an expert. He thought there might be an internal blockage of the power supplies LCW system and began repairs.

D0 experimenters made a 30-minute access to their collision hall during the B:V803 repairs.

Pbar resumed stacking at 9:14 PM. The Booster techs couldn't find a problem with B:V803. It was turned back on and operated normally.

Friday May 23
The midnight shift began with TeV expert conducting studies, with Pbar stacking, and with MiniBooNe taking beam.

At 2:09 AM, two Booster RF stations (BRF5 & 6) tripped off due to a vacuum burst. They both reset okay.

Recycler experts began taking protons for studies at 4:14 AM.

Operations put the TeV into shot setup at 6 AM.

At 7:55 AM, two Booster RF stations (BRF13 & 14) tripped off due to a vacuum burst. They both reset okay.

Today's Plan
The plan for today and this weekend through Memorial Day is to stack and store. Next week, the Fermi accelerators will be shutdown while ComEd replaces a bad high voltage power pole starting Tuesday. The job should take three days.

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 5/23/2003   email Fermilab